OpenStudy (anonymous):

What is the EXACT area of the circle with a diameter of 1.2 inch? a. 2.4 pi inches ^2 b. 1.44 pi inches ^2 c. 0.36 pi inches ^2 d. 0.18 pi inches ^2

OpenStudy (mayankdevnani):

\[\huge area ~of~circle=\pi \times r^2\]

OpenStudy (mayankdevnani):

if d= 1.2 inch , so what is the radius? r=d/2=1.2/2

OpenStudy (mayankdevnani):

so calculate radius? @bettni

OpenStudy (anonymous):

0.6

OpenStudy (mayankdevnani):

correct

OpenStudy (mayankdevnani):

so plug this value in area of circle formula

OpenStudy (anonymous):

i did already, it didnt come out

OpenStudy (mayankdevnani):

\[\huge \pi \times r^2=\pi \times 0.6 \times 0.6\] what you get when you solve it?

OpenStudy (anonymous):

1.13

OpenStudy (mayankdevnani):

no.. you are wrong

OpenStudy (mayankdevnani):

\[\huge 0.6 \times 0.6=0.36\]